OpenStudy (anonymous):

There are (4^2)^3 ⋅ 4^0 horses on a stud farm. What is the total number of horses on the farm?

OpenStudy (anonymous):

@nnesha

Nnesha (nnesha):

when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

Nnesha (nnesha):

ohh forgot the exponent rule \[\huge\rm (x^m)^n =x^{m \times n }\] first apply this

OpenStudy (anonymous):

@nnesha so the answer is 4^6?

Nnesha (nnesha):

yes right!

OpenStudy (anonymous):

@nnesha can you help with one more?

Nnesha (nnesha):

i'll try

OpenStudy (anonymous):

What is the value of 3 to the power 2 over 3 to the power 4?

Nnesha (nnesha):

and when we divide same base , `subtract` their exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

OpenStudy (anonymous):

options are : 1/81, 1/27, 1/9, 1/3

Nnesha (nnesha):

for exxponents you should remember these exponent rules

OpenStudy (anonymous):

right that would be 3^-2 but that's not an option

Nnesha (nnesha):

yes now we should convert negative to positive exponent \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\]

Nnesha (nnesha):

flip the fraction ) when u flip it the sign of the exponent would change

OpenStudy (anonymous):

so the answer is 1/3??

Nnesha (nnesha):

what about the exponent ?

OpenStudy (anonymous):

none of the options have exponents

Nnesha (nnesha):

oh well at the end we will not have the exponent ethr \[3^{-2} =???\]

OpenStudy (anonymous):

im confused.. is the answer not 1/3/?

Nnesha (nnesha):

no

Nnesha (nnesha):

look at this exponent rule when i flipped the fraction i just changed the sign exponent would stay the same \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\]